# Very Low Complexity Speech Synthesis Using Framewise Autoregressive GAN (FARGAN) with Pitch Prediction

### v1

@@ Abstract
neural vocoder, speech synthesis, GAN
@@ I Introduction
The aforementioned autoregressive vocoders all rely on explicit density estimation to synthesize the speech waveform through conditional sampling, leading to two limitations. First, their structure requires the use of teacher-forcing [11], which leads to a domain gap between training and inference that sometimes limits quality. Second, it prevents direct signal generation and the use of more advanced loss functions, as done by GAN [12] vocoders such as MelGAN [13] and HiFi-GAN [14]. On the other hand, according to [15], “autoregressive models possess an inductive bias towards learning pitch and phase”. The authors argue that the phase evolution of a periodic signal is analogous to the cumulative sum problem which is easier to learn for autoregressive models than for e.g. CNNs which are limited by their finite receptive field. The authors’ proposed CARGAN model uses past context as implicit pitch conditioning and is shown to more accurately represent the pitch compared to the non-autoregressive HiFi-GAN. Although CARGAN relies on teacher-forcing with respect to its autoregressive component, it can still be trained adversarially provided that the chunk size is sufficiently large.
@@ II FARGAN Overview
The subframe network shown in Fig. 2 is responsible for the autoregressive property of FARGAN. The signal produced at the subframe network’s output feeds back into its input in two ways. First, the last generated subframe is directly fed back to the input as a way of ensuring continuity of the signal. That autoregressive behavior is similar to that of CARGAN [15], except for the much smaller durations involved (40 samples instead of 512). The second and most important feedback path involves pitch. The input pitch period is not only used for conditioning, but is also directly used to look back in the synthesis history and extract the signal exactly one pitch period before the current subframe. For voiced speech, those samples tend to be an accurate prediction for the current subframe being synthesized. To handle cases where the pitch period $T$ may be shorter than the subframe size $N$, the predicted signal $p(n)$ is obtained by
@@
We find that simply feeding the autoregressive components back to the input of the subframe network does not result in optimal use of that information – likely due to vanishing gradient. For that reason we also feed them to all the other subframe layers. In the case of the pitch prediction, we also add a gate that avoids the prediction being used for unvoiced speech. The gate values are computed from the conditioning in the same way as the gain described above.
@@ III-B Adversarial Training
For adversarial training we use spectrogram discriminators at multiple resolutions as proposed in [18]. We follow the architecture from [16] and use 6 STFT discriminators $D_{k}$ with the modifications described in [19]: Each discriminator takes as input a log-magnitude spectrogram computed from size-$2^{k+5}$ STFTs with 75% overlap. To simplify notation, we use $D_{k}(x)$ and $D_{k}(\hat{x})$, treating the log-magnitude STFT transform of $x$ and $\hat{x}$ as part of the discriminator. We furthermore apply strides along the frequency axis to keep the frequency range of the receptive fields constant. This has been found to increase the ability of discriminators with high frequency resolution to detect inter-harmonic noise. Finally, we concatenate a two-dimensional frequency positional sine-cosine embedding to the input channels of every 2d-convolutional layer.
@@ IV Experiments & Results
We evaluated the algorithms using 192 clean English speech clips from the NTT Multi-Lingual Speech Database for Telephonometry and 192 clean English clips from the PTDB-TUG [30] database [^2] Audio samples are available at https://ahmed-fau.github.io/fargan_demo/. No items from these databases were included in the training data.
We first evaluated all the vocoders objectively using PESQ [31], WARP-Q [32], as well as mean pitch accuracy (MPE), as measured in [16]. Table I shows the objective evaluation results, demonstrating that FARGAN achieves better pitch accuracy than all other vocoders. They also demonstrate the effectiveness of the autoregressive components and show that explicit pitch prediction can help achieve both higher quality and better pitch accuracy. Although PESQ and WARP-Q results are provided for all algorithms, these metrics tend to be less reliable when comparing different families of algorithms (as opposed for variants of the same base algorithm).
Figure 3: P.808 mean opinion score (MOS) results including the 95% confidence intervals.
@@ V Conclusion
We have demonstrated a very-low-complexity GAN vocoder that uses pitch-prediction-based autoregression to achieve high-quality speech synthesis using only 600 MFLOPS. The proposed FARGAN vocoder achieves both higher quality and lower complexity when compared to other existing low-complexity vocoders. Moreover, it matches the quality of state-of-the-art high-complexity vocoders. We believe the demonstrated reduction in vocoder complexity opens the way for neural vocoders being used in new applications, including low-power embedded systems.

### v2

@@ Abstract
GAN vocoder, speech synthesis, DDSP
@@ I Introduction
The aforementioned autoregressive vocoders all rely on explicit density estimation to synthesize the speech waveform through conditional sampling, leading to two limitations. First, their structure requires the use of teacher-forcing [11], which leads to a exposure bias [12], a domain gap between training and inference that sometimes limits quality. Second, it prevents direct signal generation and the use of more advanced loss functions, as done by GAN [13] vocoders such as MelGAN [14], HiFi-GAN [15], and BigVGAN [16]. On the other hand, according to [17], “autoregressive models possess an inductive bias towards learning pitch and phase”. The authors argue that the phase evolution of a periodic signal is analogous to the cumulative sum problem which is easier to learn for autoregressive models than for e.g. CNNs which are limited by their finite receptive field. The authors’ proposed CARGAN model uses past context as implicit pitch conditioning and is shown to more accurately represent the pitch compared to the non-autoregressive HiFi-GAN. Although CARGAN relies on teacher-forcing with respect to its autoregressive component, it can still be trained adversarially provided that the chunk size is sufficiently large.
@@ II FARGAN Overview
The subframe network shown in Fig. 2 is responsible for the autoregressive property of FARGAN. The signal produced at the subframe network’s output feeds back into its input in two ways. First, the last generated subframe is directly fed back to the input as a way of ensuring continuity of the signal. That autoregressive behavior is similar to that of CARGAN [17], except for the much smaller durations involved (40 samples instead of 512). The second and most important feedback path involves pitch. The input pitch period is not only used for conditioning, but is also directly used to look back in the synthesis history and extract the signal exactly one pitch period before the current subframe. For voiced speech, those samples tend to be an accurate prediction for the current subframe being synthesized. However, as a consequence of using pitch prediction directly, the proposed model cannot be easily adapted to synthesize general audio and music.
To handle cases where the pitch period $T$ is shorter than the subframe size $N$, the predicted signal $p(n)$ is obtained by
@@
We find that simply feeding the autoregressive components back to the input of the subframe network does not result in optimal use of that information – likely due to vanishing gradient. For that reason we also feed them to all the other subframe layers. Similar to skip connections, we find that although it does not significantly improve the final output quality, it helps stabilize and speed up convergence. In the case of the pitch prediction, we also add a gate that avoids the prediction being used for unvoiced speech. The gate values are computed from the conditioning in the same way as the gain described above.
@@ III-B Adversarial Training
For adversarial training, we use multi-resolution magnitude-STFT discriminators similar to [20]. This choice is motivated by the observation that otherwise popular time domain-discriminators ([14, 15, 21]) for vocoder training failed to improve (and in fact even degraded) quality for two previous block-wise signal processing models, namely the FWGAN vocoder [18] and the NoLACE enhancement model [22]. In both cases, the multi-scale and the multi-period discriminators from [15] would quickly win against the generators indicating that these discriminators are capable of detecting (potentially irrelevant) irregularities in the generated signals that the generators were not able to remove to a satisfactory degree. A possible explanation for this behavior could be the block-wise signal processing itself and the fact that small temporal irregularities are easier detected from a raw time-domain signal than from a log-magnitude spectrogram.
We follow the architecture from [18] and use 6 STFT discriminators $D_{k}$ with the modifications described in [22]: Each discriminator takes as input a log-magnitude spectrogram computed from size-$2^{k+5}$ STFTs with 75% overlap. To simplify notation, we use $D_{k}(x)$ and $D_{k}(\hat{x})$, treating the log-magnitude STFT transform of $x$ and $\hat{x}$ as part of the discriminator. We furthermore apply strides along the frequency axis to keep the frequency range of the receptive fields constant. This has been found to increase the ability of discriminators with high frequency resolution to detect inter-harmonic noise. Finally, we concatenate a two-dimensional frequency positional sine-cosine embedding to the input channels of every 2d-convolutional layer.
@@ IV Experiments & Results
We evaluated the algorithms using 192 clean English speech clips from the NTT Multi-Lingual Speech Database for Telephonometry and 192 clean English clips from the PTDB-TUG [33] database. No items from these databases were included in the training data. Audio samples synthesizing clean speech, singing voice and noisy speech samples are available on a demo page  [^2] https://ahmed-fau.github.io/fargan_demo/. These demonstrate that FARGAN generalizes well to unseen conditions considering that no singing or noisy speech was used at training time.
We first evaluated all the vocoders objectively using PESQ [34], WARP-Q [35], as well as mean pitch accuracy (MPE), as measured in [18]. Table I shows the objective evaluation results, demonstrating that FARGAN achieves better pitch accuracy than all other vocoders. All three objective metrics are in agreement in demonstrating the effectiveness of the autoregressive components and show that explicit pitch prediction can help achieve both higher quality and better pitch accuracy. Although objective metrics are designed to correlate with subjective quality to a certain degree, comparing very different families of algorithms is a notoriously difficult task for these metrics. In this case, we notice that PESQ and WARP-Q give opposite rankings for FARGAN, HiFiGAN v1, and CARGAN, so a subjective evaluation is needed.
Figure 3: P.808 mean opinion score (MOS) results including the 95% confidence intervals. FARGAN large, HiFi-GAN v1 and CARGAN are statistically tied and out-perform all other vocoders with $p<0.05$.
@@ V Conclusion
We have demonstrated a very-low-complexity GAN vocoder that uses pitch-prediction-based autoregression to achieve high-quality speech synthesis using only 600 MFLOPS. The proposed FARGAN vocoder achieves both higher quality and lower complexity when compared to other existing low-complexity vocoders (LPCNet, HiFi-GAN v3, FWGAN). Moreover, it matches the quality of state-of-the-art high-complexity vocoders (HiFi-GAN v1, CARGAN). We believe the demonstrated reduction in vocoder complexity opens the way for neural vocoders being used in new applications, including low-power embedded systems.